# Optimal Computation in Leaderless and Multi-Leader Disconnected Anonymous Dynamic Networks

### 20 Feb 2024 v5

@@ 20         Return $(t,S)$
Let $\mathcal{V}$ be the view of a process in a $T$-union-connected network of size $n$ taken at round $t^{\prime}$, and let Listing 1 return $(t,S)$ on input $\mathcal{V}$. Assume that one of the following conditions holds:
@@
    $t^{\prime}\geq 2Tn$.
Then, $0\leq t\leq Tn$, and $S$ is a homogeneous system of $k-1$ independent linear equations (with integer coefficients) in $k=|L_{t}|$ variables $x_{1}$, $x_{2}$, …, $x_{k}$. Moreover, $S$ is satisfied by assigning to $x_{i}$ the anonymity of the $i$th node of $L_{t}$, for all $1\leq i\leq k$.
@@
It is well known that, if $T=1$, information takes less than $n$ rounds to travel from a process to any other process [36]. Thus, if $T\geq 1$, it takes less than $Tn$ rounds (cf. Proposition 2.3). Since $\mathcal{V}$ is a view taken at round $2Tn$ (or after), all levels of $\mathcal{V}$ up to $L_{t^{\prime}-Tn+1}$ are complete, i.e., all nodes in the first $t^{\prime}-Tn+1$ levels of the network’s history tree also appear in the view $\mathcal{V}$.
@@
Assume Condition 2 first. Since $t^{\prime}\geq 2Tn$, all levels of $\mathcal{V}$ up to $L_{Tn+1}$ are complete. Since the anonymity of the root of $\mathcal{V}$ is $n$, there must be less than $n$ branching nodes in $\mathcal{V}$. Therefore, the first $Tn$ levels contain an interval of at least $T$ consecutive levels, say from $L_{r}$ to $L_{r+T-1}$, where all nodes are non-branching and can be partitioned into $|L_{r}|=|L_{r+T-1}|$ strands $P_{i}$.
@@
Note that a link between two processes at any round $r^{\prime}$ in the interval $[r+1,r+T]$ determines a pair of exposed nodes in $L_{r^{\prime}-1}$. Thus, by definition of $T$-union-connected network, the graph of exposed strands between $L_{r}$ and $L_{r+T-1}$ (constructed as $G$ in Line 12) is connected. It follows that the execution of Listing 1 terminates at Line 20 (as opposed to Line 6), at the latest when $t=r+T-1$. Thus, the procedure returns a pair $(t,S)$ with $0\leq t\leq r+T-1\leq Tn$. In particular, all levels of $\mathcal{V}$ up to $L_{t+1}$ are complete.
@@
Now assume Condition 1. Since $t^{\prime}\geq t+Tn$, all levels of $\mathcal{V}$ up to $L_{t+1}$ are complete in this case, as well. Since $t\geq 0$ by assumption, the execution of Listing 1 terminates at Line 20. The termination condition is met when long-enough strands are found; as proved above, this must happen while $t\leq Tn$.
@@
We have proved that, in both cases, the inequalities $0\leq t\leq Tn$ hold, and all levels of $\mathcal{V}$ up to $L_{t+1}$ are complete. Let us now examine the linear system $S$. Observe that $S$ is homogeneous because it consists of homogeneous linear equations (cf. Line 19). Also, since the spanning tree $G^{\prime}$ constructed in Line 14 has $k-1$ edges, $S$ contains $k-1$ equations. We will prove that they are linearly independent by induction on $k$. If $k=1$, there is nothing to prove. Otherwise, let $P_{i}$ be a leaf of $G^{\prime}$, and let $\{P_{i},P_{j}\}$ be its incident edge. Then, $S$ contains an equation $Q$ of the form $m_{1}x_{i}=m_{2}x_{j}$ with $m_{1}m_{2}\neq 0$. Let $S^{\prime}$ be the system obtained by removing $Q$ from $S$; equivalently, $S^{\prime}$ corresponds to the tree obtained by removing the leaf $P_{i}$ from $G^{\prime}$. By the inductive hypothesis, no linear combination of equations in $S^{\prime}$ yields $0=0$. On the other hand, if $Q$ is involved in a linear combination with a non-zero coefficient, then the variable $x_{i}$ cannot vanish, because it only appears in $Q$. Therefore, the equations in $S$ are independent.
@@
It remains to prove that a solution to $S$ is given by the anonymities of the nodes of $L_{t}$. It was shown in [24, Lemma 4.1] that, if $v_{1}$ and $v_{2}$ are exposed in $\mathcal{V}$, as well as in the history tree containing $\mathcal{V}$, with multiplicity $(m_{1},m_{2})$, then $m_{1}a(v_{1})=m_{2}a(v_{2})$. [^17] If $v_{1}$ and $v_{2}$ are exposed in $\mathcal{V}$ but not in the underlying history tree, then they have some children not in $\mathcal{V}$, and therefore the equation may not hold. To conclude our proof, it is sufficient to note that, since the nodes of a strand $P_{i}$ are non-branching in $\mathcal{V}$ as well as in the underlying history tree (recall that all levels of $\mathcal{V}$ up to $L_{t+1}$ are complete), they all have the same anonymity, which is the anonymity of the ending node $w_{i}\in L_{t}$. ∎
@@
There is an algorithm that computes $F_{R}$ in all $T$-union-connected anonymous networks with no leader and stabilizes in at most $2Tn$ rounds, assuming no knowledge of $T$ or $n$.
@@
Our local algorithm is as follows. Run Listing 1 on the process’ view $\mathcal{V}$, obtaining a pair $(t,S)$. If $t=-1$ or $S$ is not a homogeneous system of $k-1$ independent linear equations in $k$ variables, output “Unknown”. Otherwise, since the rank of the coefficient matrix of $S$ is $k-1$, the general solution to $S$ has exactly one free parameter, due to the Rouché–Capelli theorem. Therefore, by Gaussian elimination, it is possible to express every variable $x_{i}$ as a rational multiple of $x_{1}$, i.e., $x_{i}=\alpha_{i}x_{1}$ for some $\alpha_{i}\in\mathbb{Q}^{+}$ (note that the coefficients of $S$ are integers). Let $L_{t}=\{w_{1},w_{2},\dots,w_{k}\}$ and $L_{0}=\{v_{1},v_{2},\dots,v_{k^{\prime}}\}$. For every node $v_{i}\in L_{0}$, define $\beta_{i}\in\mathbb{Q}^{+}$ as $\beta_{i}=\sum_{w_{j}\in L_{t}\text{ descendant of }v_{i}}\alpha_{j}$, and let $\beta=\sum_{i}\beta_{i}$. Then, output
@@
The correctness and stabilization time of the above algorithm directly follow from Lemma 3.1. Specifically, at any round $t^{\prime}\geq 2Tn$, Condition 2 of Lemma 3.1 is met, and the system $S$ is satisfied by the anonymities of the nodes in $L_{t}$. Thus, $a(v_{i})=\alpha_{i}a(v_{1})$ for all $v_{i}\in L_{0}$, and therefore $\beta_{i}/\beta=a(v_{i})/n$. We conclude that, for any input assignment $\lambda$, the algorithm stabilizes on the correct output $\frac{1}{n}\cdot\mu_{\lambda}$ within $2Tn$ rounds. ∎
3.2 Terminating Algorithm
We will now give a certificate of correctness that can be used to turn the stabilizing algorithm of Theorem 3.2 into a terminating algorithm. The certificate relies on a-priori knowledge of the dynamic disconnectivity $T$ and an upper bound $N$ on the size of the network $n$; these assumptions are justified by Proposition 2.2 and Proposition 5.3, respectively.
@@ 3.2 Terminating Algorithm
There is an algorithm that computes $F_{R}$ in all $T$-union-connected anonymous networks with no leader and terminates in at most $T(n+N)$ rounds, assuming that $T$ and an upper bound $N\geq n$ are known to all processes. [^18] If the dynamic diameter $D$ of the network is known, the termination time improves to $Tn+D$ rounds.
@@
Our terminating algorithm is as follows. Run Listing 1 on the process’ view $\mathcal{V}$, obtaining a pair $(t,S)$, and then do the same computations as in the algorithm in Theorem 3.2. If $t\geq 0$ and the current round $t^{\prime}$ satisfies $t^{\prime}\geq t+TN$, then the output is correct, and the process terminates.
The correctness of this algorithm is a direct consequence of Lemma 3.1. Indeed, the algorithm only terminates when $t\geq 0$ and $t^{\prime}\geq t+TN\geq t+Tn$, and hence it gives the correct output because Condition 1 of Lemma 3.1 is met. As for the running time, assume that $t^{\prime}=T(n+N)\geq 2Tn$. Since Condition 2 of Lemma 3.1 is met, we have $0\leq t\leq Tn$. Thus, $t^{\prime}=T(n+N)\geq t+Tn$, and the algorithm terminates at round $t^{\prime}$. ∎
4 Computation in Networks with Leaders
We will give a stabilizing and a terminating algorithm that efficiently compute the Generalized Counting function $F_{GC}$ in all networks with $\ell\geq 1$ leaders and finite dynamic disconnectivity $T$. Therefore, all multi-aggregate functions are efficiently computable as well, due to [24, Theorem 2.1]. Moreover, Proposition 5.4 states that no other functions are computable in networks with leaders, and Proposition 5.6 shows that our algorithms are asymptotically optimal for any fixed $\ell\geq 1$.
4.1 Stabilizing Algorithm
We will once again make use of the subroutine in Listing 1, this time assuming that the number of leaders $\ell\geq 1$ is known to all processes. This assumption is justified by Proposition 5.5.
@@ 4.1 Stabilizing Algorithm
There is an algorithm that computes $F_{GC}$ in all $T$-union-connected anonymous networks with $\ell\geq 1$ leaders and stabilizes in at most $2Tn$ rounds, assuming that $\ell$ is known to all processes, but assuming no knowledge of $T$ or $n$.
@@
The algorithm proceeds as in Theorem 3.2. When the fractions $\beta_{1}$, $\beta_{2}$, …, $\beta_{k^{\prime}}$ have been computed, as well as their sum $\beta$, we perform the following additional steps. Let $L_{0}=\{v_{1},v_{2},\dots,v_{k^{\prime}}\}$, and let $\{v_{j_{1}},v_{j_{2}},\dots,v_{j_{l}}\}\subseteq L_{0}$ be the set of nodes in $L_{0}$ representing leader processes, [^19] In general we have $l\leq\ell$, because some nodes of $L_{0}$ may represent more than one leader. i.e., such that $\text{label}(v_{j_{i}})$ has the leader flag set for all $1\leq i\leq l$. Compute $\beta^{\prime}=\sum_{i=1}^{l}\beta_{j_{i}}$ and $\gamma_{i}=\ell\beta_{i}/\beta^{\prime}$ for all $1\leq i\leq k^{\prime}$, and output
@@
The correctness follows from the fact that, as shown in Theorem 3.2, at any round $\geq 2Tn$ we have $\beta_{i}/\beta=a(v_{i})/n$ for all $1\leq i\leq k^{\prime}$. Adding up these equations for all $i\in\{j_{1},j_{2},\dots,j_{l}\}$, we obtain $\beta^{\prime}/\beta=\ell/n$, and therefore $n=\ell\beta/\beta^{\prime}$. We conclude that
@@
Thus, within $2Tn$ rounds, the algorithm stably outputs the anonymities of all nodes in $L_{0}$. As observed in Section 2, this is equivalent to computing the Generalized Counting function $F_{GC}$. ∎
4.2 Terminating Algorithm
We will now present the main result of this paper. As already remarked, giving an efficient certificate of correctness for the (Generalized) Counting problem with multiple leaders is a highly non-trivial task for which a radically different approach is required. Note that it is not possible to simply adapt the single-leader algorithm in [24] by setting the anonymity of the leader node in the history tree to $\ell$ instead of $1$. Indeed, as soon as some leaders get disambiguated, the leader node splits into several children nodes whose anonymities are unknown (we only know that their sum is $\ell$). There is no way around this difficulty other than developing a new technique.

### 05 Apr 2024 v6

@@ 20         Return $(t,S)$
Let $\mathcal{V}$ be the view of a process in a $\tau$-union-connected network of size $n$ taken at round $t^{\prime}$, and let Listing 1 return $(t,S)$ on input $\mathcal{V}$. Assume that one of the following conditions holds:
@@
    $t^{\prime}\geq\tau(2n-c-1)+1$, where $c$ is the number of distinct inputs held by processes at round 00.
Then, $0\leq t\leq\tau(n-c)$, and $S$ is a homogeneous system of $k-1$ independent linear equations (with integer coefficients) in $k=|L_{t}|$ variables $x_{1}$, $x_{2}$, …, $x_{k}$. Moreover, $S$ is satisfied by assigning to $x_{i}$ the anonymity of the $i$th node of $L_{t}$, for all $1\leq i\leq k$.
@@
If $\tau=1$, it takes at most $n-1$ rounds for information to travel from a process to any other process, due to Corollary 3.4. In general, if $\tau\geq 1$, it takes at most $\tau(n-1)$ rounds, by Proposition 2.4. Therefore, since $\mathcal{V}$ is a view taken at round $t^{\prime}$, all levels of $\mathcal{V}$ up to $L_{t^{\prime}-\tau(n-1)}$ are complete (recall that a level of a view is complete if it has no missing nodes).
@@
Assume Condition 2 first. Since $t^{\prime}\geq\tau(2n-c-1)+1$, all levels of $\mathcal{V}$ up to $L_{\tau(n-c)+1}$ are complete. Thus, level $L_{0}$ is complete, and therefore it has exactly $c$ nodes, because two processes are distinguishable at round 00 if and only if they have distinct inputs. Since the sum of the anonymities of these $c$ nodes is $n$, there must be fewer than $n-c$ branching nodes in $\mathcal{V}$ (excluding the root). Hence, the levels up to $L_{\tau(n-c)}$ contain an interval of at least $\tau$ consecutive levels, say from $L_{r}$ to $L_{r+\tau-1}$, where all nodes are non-branching and can be partitioned into $|L_{r}|=|L_{r+\tau-1}|$ strands $P_{1}$, $P_{2}$, …, $P_{|L_{r}|}$.
@@
Note that a link between two processes at any round $r^{\prime}$ in the interval $[r+1,r+\tau]$ determines a pair of exposed nodes in $L_{r^{\prime}-1}$. Thus, by definition of $\tau$-union-connected network, the graph of exposed strands between $L_{r}$ and $L_{r+\tau-1}$ (constructed as $G$ in Line 12) is connected. It follows that the execution of Listing 1 terminates at Line 20 (as opposed to Line 6) whenever $t\geq r+\tau-1$. Thus, the procedure returns a pair $(t,S)$ with $0\leq t\leq r+\tau-1\leq\tau(n-c)$. In particular, all levels of $\mathcal{V}$ up to $L_{t+1}$ are complete.
@@
Now assume Condition 1. Since $t^{\prime}\geq t+\tau(n-1)+1$, all levels of $\mathcal{V}$ up to $L_{t+1}$ are complete in this case, as well. Since $t\geq 0$ by assumption, the execution of Listing 1 terminates at Line 20. The termination condition is met when long-enough strands are found; as proved above, this event must occur while $t\leq\tau(n-c)$.
@@
We have proved that, in both cases, the inequalities $0\leq t\leq\tau(n-c)$ hold, and all levels of $\mathcal{V}$ up to $L_{t+1}$ are complete. Let us now examine the linear system $S$. Observe that $S$ is homogeneous because it consists of homogeneous linear equations (cf. Line 19). Also, since the spanning tree $G^{\prime}$ constructed at Line 14 has $k-1$ edges, $S$ contains $k-1$ equations. We will prove that they are linearly independent by induction on $k$. If $k=1$, there is nothing to prove. Otherwise, let $P_{i}$ be a leaf of $G^{\prime}$, and let $\{P_{i},P_{j}\}$ be its incident edge. Then, $S$ contains an equation $Q$ of the form $m_{1}x_{i}=m_{2}x_{j}$ with $m_{1}m_{2}\neq 0$. Let $S^{\prime}$ be the system obtained by removing $Q$ from $S$; equivalently, $S^{\prime}$ corresponds to the tree obtained by removing the leaf $P_{i}$ from $G^{\prime}$. By the inductive hypothesis, no linear combination of equations in $S^{\prime}$ yields $0=0$. On the other hand, if $Q$ is involved in a linear combination with a non-zero coefficient, then the variable $x_{i}$ cannot vanish, because it only appears in $Q$. Therefore, the equations in $S$ are independent.
@@
It remains to prove that a solution to $S$ is given by the anonymities of the nodes of $L_{t}$. Due to Lemma 4.1, if $v_{1}$ and $v_{2}$ are exposed in $\mathcal{V}$, as well as in the history tree containing $\mathcal{V}$, with multiplicity $(m_{1},m_{2})$, then $m_{1}a(v_{1})=m_{2}a(v_{2})$. To conclude our proof, it is sufficient to note that, since the nodes of a strand $P_{i}$ are non-branching in $\mathcal{V}$ as well as in the underlying history tree (recall that all levels of $\mathcal{V}$ up to $L_{t+1}$ are complete), they all have the same anonymity, which is the anonymity of the ending node $w_{i}\in L_{t}$. ∎
Diameter bounds. We can also restate Lemma 4.2 with respect to the dynamic diameter $d$ of the network.
Corollary 4.3.
The statement of Lemma 4.2 remains valid if the two conditions are replaced with:
  1. 1.
    $t\geq 0$ and $t^{\prime}\geq t+d+1$, or
  2. 2.
    $t^{\prime}\geq\tau(n-c)+d+1$, where $c$ is the number of distinct inputs held by processes at round 00,
where $d$ is the dynamic diameter of the network (or an upper bound thereof).
Proof.
Note that it takes at most $d$ rounds for information to travel across the network. Hence, the proof of Lemma 4.2 can be repeated verbatim, substituting the term $\tau(n-1)$ with $d$ throughout. ∎
4.2 Stabilizing Algorithm for Leaderless Networks
As a first application of Lemma 4.2, we will give stabilizing algorithm that efficiently computes the Input Frequency function $F_{IF}$ in all leaderless networks with a finite dynamic disconnectivity $\tau$. As a consequence, all frequency-based functions are efficiently computable as well, due to Proposition 2.2. Moreover, Proposition 6.3 states that no other functions are computable in leaderless networks, and Proposition 6.9 shows that our algorithm is asymptotically optimal.
@@ 4.2 Stabilizing Algorithm for Leaderless Networks
There is an algorithm that computes $F_{IF}$ in all $\tau$-union-connected anonymous networks with no leader and stabilizes in at most $\tau(2n-2)+1$ rounds, assuming no knowledge of $\tau$ or $n$.
@@
Our local algorithm is as follows. Run Listing 1 on the process’ view $\mathcal{V}$, obtaining a pair $(t,S)$. If $t=-1$ or $S$ is not a homogeneous system of $k-1$ independent linear equations in $k$ variables, output “Unknown”. Otherwise, since the rank of the coefficient matrix of $S$ is $k-1$, the general solution to $S$ has exactly one free parameter, due to the Rouché–Capelli theorem. Therefore, by Gaussian elimination, it is possible to express every variable $x_{i}$ as a rational multiple of $x_{1}$, i.e., $x_{i}=\alpha_{i}x_{1}$ for some $\alpha_{i}\in\mathbb{Q}^{+}$ (recall that the coefficients of $S$ are integers). Let $L_{t}=\{w_{1},w_{2},\dots,w_{k}\}$ and $L_{0}=\{v_{1},v_{2},\dots,v_{k^{\prime}}\}$. For every node $v_{i}\in L_{0}$, define $\beta_{i}\in\mathbb{Q}^{+}$ as $\beta_{i}=\sum_{w_{j}\in L_{t}\text{ descendant of }v_{i}}\alpha_{j}$, and let $\beta=\sum_{i}\beta_{i}$. Then, output
@@
The correctness and stabilization time of the above algorithm directly follow from Lemma 4.2. Specifically, at any round $t^{\prime}\geq\tau(2n-2)+1$, Condition 2 of Lemma 4.2 is met (since $c\geq 1$), and the system $S$ is satisfied by the anonymities of the nodes in $L_{t}$. Thus, $a(v_{i})=\alpha_{i}a(v_{1})$ for all $v_{i}\in L_{0}$, and therefore $\beta_{i}/\beta=a(v_{i})/n$. We conclude that, for any input assignment $\lambda$, the algorithm stabilizes on the correct output $\frac{1}{n}\cdot\mu_{\lambda}$ within $2\tau n$ rounds. ∎
4.3 Terminating Algorithm for Leaderless Networks
We will now give a certificate of correctness that can be used to turn the stabilizing algorithm of Section 4.2 into a terminating algorithm. The certificate relies on a-priori knowledge of the dynamic disconnectivity $\tau$ and an upper bound $N$ on the size of the network $n$; these assumptions are justified by Proposition 2.3 and Proposition 6.4, respectively. Again, Proposition 6.9 shows that our algorithm is asymptotically optimal.
@@ 4.3 Terminating Algorithm for Leaderless Networks
There is an algorithm that computes $F_{IF}$ in all $\tau$-union-connected anonymous networks with no leader and terminates in at most $\tau(n+N-2)+1$ rounds, assuming that $\tau$ and an upper bound $N\geq n$ are known to all processes.
@@
Assume that $\tau$ and $N$ are known. Our terminating local algorithm is as follows. Run Listing 1 on the process’ view $\mathcal{V}$, obtaining a pair $(t,S)$, and then do the same computations as in the algorithm for Theorem 4.4. If $t\geq 0$ and the current round $t^{\prime}$ satisfies $t^{\prime}\geq t+\tau(N-1)+1$, then the output is correct, and the process terminates.
The correctness of this algorithm is a direct consequence of Lemma 4.2. Indeed, the algorithm only terminates when $t\geq 0$ and $t^{\prime}\geq t+\tau(N-1)+1\geq t+\tau(n-1)+1$, and hence it gives the correct output because Condition 1 of Lemma 4.2 is met. As for the running time, assume that $t^{\prime}=\tau(n+N-2)+1\geq\tau(2n-c-1)+1$. Since Condition 2 of Lemma 4.2 is met, we have $0\leq t\leq\tau(n-c)\leq\tau(n-1)$. Thus, $t^{\prime}=\tau(n+N-2)+1\geq t+\tau(N-1)+1$, and the algorithm terminates at round $t^{\prime}$. ∎
We can also trade the knowledge of $\tau$ and $N$ for the knowledge of the dynamic diameter $d$ of the network.
Corollary 4.6.
There is an algorithm that computes $F_{IF}$ in all $\tau$-union-connected anonymous networks with no leader and terminates in at most $\tau(n-1)+d+1\leq dn+1$ rounds, assuming that the dynamic diameter (or an upper bound thereof) $d$ is known to all processes, but assuming no knowledge of $\tau$ or $n$.
Proof.
We can repeat the proof of Theorem 4.5 verbatim, replacing the termination condition with $t^{\prime}\geq t+d+1$ and invoking Corollary 4.3 in lieu of Lemma 4.2. ∎
4.4 Stabilizing Algorithm for Networks with Leaders
To conclude this section, we will give a stabilizing algorithm that efficiently computes the Input Multiset function $F_{IM}$ in all networks with $\ell\geq 1$ leaders and a finite dynamic disconnectivity $\tau$. Therefore, all multiset-based functions are efficiently computable as well, due to Proposition 2.1. Moreover, Proposition 6.1 states that no other functions are computable in networks with leaders, and Proposition 6.8 shows that our algorithm is asymptotically optimal. We will once again make use of the subroutine in Listing 1, this time assuming that the number of leaders $\ell\geq 1$ is known to all processes. This assumption is justified by Proposition 6.2.
@@ 4.4 Stabilizing Algorithm for Networks with Leaders
There is an algorithm that computes $F_{IM}$ in all $\tau$-union-connected anonymous networks with $\ell\geq 1$ leaders and stabilizes in at most $\tau(2n-3)+1$ rounds, assuming that $\ell$ is known to all processes, but assuming no knowledge of $\tau$ or $n$.
@@
The algorithm proceeds as in Theorem 4.4, with two differences. First, instead of returning “Unknown”, the new algorithm returns $\ell$. Second, when the fractions $\beta_{1}$, $\beta_{2}$, …, $\beta_{k^{\prime}}$ have been computed, as well as their sum $\beta$, we perform the following additional steps. Let $L_{0}=\{v_{1},v_{2},\dots,v_{k^{\prime}}\}$, and let $\{v_{j_{1}},v_{j_{2}},\dots,v_{j_{l}}\}\subseteq L_{0}$ be the set of nodes in $L_{0}$ representing leader processes, [^22] In general we have $l\leq\ell$, because some nodes of $L_{0}$ may represent more than one leader. i.e., such that $\text{label}(v_{j_{i}})$ has the leader flag set for all $1\leq i\leq l$. Compute $\beta^{\prime}=\sum_{i=1}^{l}\beta_{j_{i}}$ and $\gamma_{i}=\ell\beta_{i}/\beta^{\prime}$ for all $1\leq i\leq k^{\prime}$, and output
@@
The correctness follows from the fact that, as shown in Theorem 4.4, at any round $t^{\prime}\geq\tau(2n-c-1)+1$ we have $\beta_{i}/\beta=a(v_{i})/n$ for all $1\leq i\leq k^{\prime}$. Adding up these equations for all $i\in\{j_{1},j_{2},\dots,j_{l}\}$, we obtain $\beta^{\prime}/\beta=\ell/n$, and therefore $n=\ell\beta/\beta^{\prime}$. We conclude that
@@
Thus, within $\tau(2n-c-1)+1$ rounds, the algorithm stably outputs the anonymities of all nodes in $L_{0}$. As observed in Section 2, this is equivalent to computing the Input Multiset function $F_{IM}$.
If there are both leader and non-leader processes in the system, there are $c\geq 2$ distinct inputs, and therefore the stabilization time is at most $\tau(2n-3)+1$ rounds, as desired. In the special case where all processes are leaders, we have $\ell=n$, which is the algorithm’s output at every round since round 00 (recall that, by design, the algorithm returns $\ell$ instead of “Unknown”). Thus, the stabilization time is $\tau(2n-3)+1$ rounds in any case. ∎
5 Terminating Algorithm for Networks with Leaders
We will now present the main result of this paper. As already remarked, giving an efficient certificate of correctness for the Input Multiset function with one or more leaders is a highly non-trivial task for which a radically new approach is required. In fact, there are two crucial difficulties to overcome.
Firstly, the strategies developed in Section 4 are too shallow and ineffective even for networks with a unique leader, which are the easiest to treat. In Section 5.1, we will give counterexamples to some naive termination strategies that one may devise in an attempt to generalize those in Section 4. This indicates that an entirely different technique is necessary.
Secondly, networks with more than one leader significantly add to the difficulty of the problem. For instance, once a terminating algorithm for networks with a unique leader has been designed, one may be tempted to simply adapt it to the multi-leader case by setting the anonymity of the leader node in the history tree to $\ell>1$ instead of $1$. However, this approach would be inconsistent. Indeed, while the history tree of a network with a unique leader contains a single leader branch, all of whose nodes have anonymity $1$, this may not be the case for networks with multiple leaders. In such a network, as soon as some leaders get disambiguated, the leader node branches into several children nodes whose anonymities are unknown (we only know that their sum is $\ell$). Moreover, some leader nodes may be missing from the view of other leaders. Thus, in the case of multi-leader networks, we must deal with the fact that even the view of a leader may have levels where the sum of the anonymities of any subset of nodes is unknown.
Section 5.3 contains the technical core of our algorithm. Here we develop a subroutine that, in $O(\tau\ell n)$ rounds, counts the number of processes in a network assuming that its history tree contains a leader node of known anonymity whose descendants are non-branching for sufficiently many rounds. In particular, in the case of networks with a unique leader, this subroutine yields a full-fledged linear-time terminating algorithm for the Counting problem (because in this case the leader nodes are necessarily non-branching).
In Section 5.2, we approach the general multi-leader problem by repeatedly guessing the anonymity of a leader node in the history tree and invoking the subroutine of Section 5.3 to confirm our guesses. This procedure introduces a further bottleneck that causes our final algorithm to have a running time of $O(\tau\ell^{2}n)$ rounds.
5.1 Naive Termination Strategies Are Incorrect
If we were to use a technique such as the one in Section 4.4 to devise a terminating algorithm for the Counting problem (or, more generally, for computing the Input Multiset function), our attempts would be bound to fail.
Basic naive strategy. As a first example, consider the dynamic network in Figure 4. Observe that the leader of this network always receives exactly four messages from indistinguishable non-leader processes. In turn, the processes that send messages to the leader in the first four rounds are still unaware of the processes labeled $p_{3}$. As a result, the view of the leader up to round $4$ consists of only two strands whose nodes are exposed with multiplicity $(4,1)$ at every level. Thus, according to the algorithm in Section 4.4, the leader assigns an anonymity of $4$ to all the non-leader nodes in its view, concluding that there are only five processes in the network. Essentially, for the first four rounds, the leader cannot distinguish this network from a static star graph with the leader at the center.
Figure 4: An example of a dynamic network where the naive technique of Section 4.4 fails to provide a correct termination condition. The white nodes in the history tree are not in the history of the leader at the last round; the red edges not in the view are not drawn. Same-colored processes have equal inputs. For the first four rounds, from the leader’s perspective, this network is indistinguishable from the complete bipartite graph $K_{1,4}$. Thus, throughout this time, the leader is unaware of the processes labeled $p_{3}$, and therefore cannot compute the total number of processes.
It is straightforward to generalize this example to networks with $n=k_{1}+k_{2}+1$ processes, $k_{1}+1$ of which are counted by the leader, while the other $k_{2}$ are not discovered by the leader until round $k_{1}+1$ (such as the processes labeled $p_{3}$ in Figure 4). In these networks, our naive algorithm consistently returns $k_{1}+1$ for several rounds, which can be made arbitrarily far from the true value $k_{1}+k_{2}+1$ by increasing $k_{2}$ indefinitely.
Improved naive strategy. One may conjecture that a good termination certificate would be to compute the number of processes $n^{\prime}$ according to the algorithm in Section 4.4, and then wait for a number of rounds depending on $n^{\prime}$ to confirm that no relevant nodes were missing from the view. In fact, in the previous example, simply waiting for $n^{\prime}$ rounds would suffice.
Figure 5: A dynamic network where, after level $L_{1}$, all levels in the leader’s view are identical for an arbitrarily long sequence of rounds (depending on the parameter $k$). The color schemes and stylistic conventions are as in Figure 4.
Unfortunately, this strategy fails on the network in Figure 5. Here the leader’s view at levels $L_{0}$ and $L_{1}$ causes the algorithm to count only $n^{\prime}=5$ processes (i.e., one leader, two processes represented by the purple node, and one process represented by the yellow node). Afterwards, the leader has to wait until round $k-5$ for the appearance of the node that was missing from $L_{1}$. Since $k$ is arbitrary, this type of strategy is bound to fail no matter what the waiting time is.
Challenges. Essentially, the recurring issue is that a process seems to have no way of knowing whether any level in its view of the history tree is complete; thus, it may end up terminating too soon with an incorrect output. To establish a correct termination condition, we will have to considerably develop the theory of history trees. This will be undertaken in Sections 5.2 and 5.3.
5.2 Main Algorithm